OpenStudy (anonymous):

Solve: 3(x + 4) - 4(x + 3) = 7x + 16 @Emotions and Life…

OpenStudy (anonymous):

Can you help me FoolforMath?

OpenStudy (anonymous):

3(x + 4) - 4(x + 3) = 7x + 16 3x + 12 - 4x - 12 = 7x + 16 -x = 7x + 16 -7x -7x -8x = 16 /-8 /-8 x = -2